OpenStudy (anonymous):

prove the identity: [\left( \cos2x \over \sin x \right) = \left( \cot^2x-1 \over \csc x \right)\]

OpenStudy (anonymous):

\[\left( \cos2x \over \sin x \right) = \left( \cot^2x-1 \over \csc x \right)\]

OpenStudy (mertsj):